# Lightweight Embedded FPGA Deployment of Learned Image Compression with Knowledge Distillation and Hybrid Quantization

### v3

@@ III-B Efficient GDN design for FPGA implementation
Other works [20, 21, 22] replace the GDN layer with ReLU activation to simplify the implementation. For our FPGA-based LIC architecture, we are the first to propose a custom Generalized Divisive Normalization (GDN) hardware engine to optimize RD performance. The GDN layer is essential for reducing statistical dependencies between activations in LIC models, which improves compression quality. Given an input feature map $\mathbf{x}\in\mathbb{R}^{H\times W\times C}$, where $H$, $W$, and $C$ are the height, width, and number of channels respectively, the GDN operation is defined as:
@@ III-D Mixed precision quantization
The parameters that survived pruning are then quantized towards integers arithmetic operations over FPGA. In most hardware LIC implementations, all parameters are pruned over 8-bit precision to minimize latency. Conversely, we proposed a mixed-precision quantization scheme where different model components are assigned different bit numbers. Specifically, the Generalized Divisive Normalization (GDN) layer is assigned a higher precision of 32-bit, while the rest of the model operates at 8-bit precision. This decision is motivated by the fact that the GDN layer, while only responsible for approximately 4% of the total FLOPS in the model, plays a critical role in improving the RD efficiency [31]. We recall that GDN’s primary function is to normalize activations across channels, promoting decorrelation in the latent space and improving RD efficiency. Lower precision (8-bit) introduces quantization noise in GDN’s operations such as division and normalization, which diminishes the accuracy of these transformations. By assigning 32-bit precision to the GDN layer, we reduce this quantization noise, leading to better normalization and decorrelation, ultimately enhancing rate-distortion performance, as we experimentally show also later on. Fig.6 illustrates the overall quantization strategy. Preliminarily, min and max values are extracted from the different model layers, allowing the quantizer to derive the optimal scale and zero point from the float model using the calibration subset. This subset includes 1000 random images from CLIC [32], DIV2K [33], and Open Images [34], used only for a forward pass. Symmetric quantization is achieved by setting the zero point to zero, with saturation applied to minimize quantization differences. As quantization impairs the model performance, the model undergoeas a Quantization Aware Tuning (QAT) is then tuned again for 1000 iterations over a subset of 5000 images from CLIC, DIV2K, and Open Images. Simulated quantization operations are introduced on both weights and activations, except for the GDN layer, which maintains 32-bit precision to avoid losing the benefits of accurate normalization. During each forward pass, operations use the quantized weights and activations, and gradients are computed with respect to the quantized parameters using a Straight-Through Estimator (STE) [36] to approximate gradients through the quantization process. The outcome of the quantization process is an LIC model that exclusively relies on integer operations, meeting the requirements for efficient FPGA deployment.
@@ III-E Design Space Exploration
In this section, we finally describe the hardware configuration we use for implementing the quantized LIC model model on FPGA using the Vitis AI platform [29]. Namely, we target the Xilinx ZCU102 FPGA that houses three DPU cores running at 300 MHz. We propose below a fully pipelined configuration employing the common B4096-DPU core configuration. Such configuration provides a peak of 4096 Multiply-Accumulate (MAC) operations per cycle featuring 16 input channels, 16 output channels, and processing 8 pixels per cycle. Namely, we seek a fully pipelined configuration where each of the three DPU cores operates in parallel, as depicted in Fig.7. By dividing the image into three patches sized 256 × 256 patches and processing multiple patches concurrently, the fully pipelined configuration optimizes resource utilization and minimizes end-to-end latency. Each DPU core handles a distinct patch, ensuring continuous processing and efficient use of the available computational resources. This setup dedicates each DPU to processing different image patches simultaneously, thereby maximizing throughput and reducing idle times.
@@
The model is compiled into target-specific DPU instructions in the form of an xmodel. The XIR compiler constructs an internal computation graph as an intermediate representation (IR), it then performs multiple optimizations, such as computation node fusion, and efficient instruction scheduling by exploiting inherent parallelism or exploiting data reuse on the specific target. This unified IR is broken down into subgraphs based on the corresponding control flow, for our models multiple DPU subgraphs are compiled as the custom GDN layers require separate graphs. Finally, an instruction stream is generated for each subgraph and everything is serialized into a target-specific, compiled xmodel file. This xmodel is encrypted and then deployed on the FPGA using the runtime API through the embedded processor. We deploy the models on the Zynq-UltraScale+ MPSoC DPU architecture -DPUCZDX8L, seen Fig.8. The nomenclature refers to the target FPGA, quantization bit-width and design objective, with L being low latency. The DPU provides multiple configurations to trade off PL resources for performance. The B4096 core configuration yields the highest throughput with a peak of 18×16×16×2=4096 MAC/cycle. Three B4096-DPUCZDX8L cores can are deployed on the ZCU102.
@@ III-F Analytical Model for Estimating FPS
$T_{\text{compute}}=\frac{N_{\text{workload}}}{\text{Peak Ops per Cycle}\times\eta}$
@@
The **Peak Ops per Cycle** is determined by the DPU configuration, which includes the number of input and output channels per layer, pixel parallelism, and core utilization. For each convolutional layer, the total number of operations per cycle per core is given by:
@@
For the B4096 configuration, where $PP=8$, $ICP=16$, and $OCP=16$, the peak operations per cycle per core is:
[…]
Given three DPU cores operating in parallel, the total peak operations per cycle across all cores is:
[…]
@@
$T_{\text{frame}}=\frac{N_{\text{workload}}}{\text{Peak Ops per Cycle (total)}\times\eta}$
@@
FPS=Peak Ops per Cycle ($\text{FPS}=\frac{\text{Peak Ops per Cycle (total)}\times\eta}{N_{\text{workload}}}$
@@ IV Experiments and Results
IV-A Training the teacher and distilling the students
To start with, we train from scratch the LIC model [3] for validating our experimental setup using the setup described in the original paper. The model is trained using a randomized train/validate/test dataset split of the combined CLIC, DIV2K, a subset of the Open Images datasets over batches of 256x256 randomly cropped patches and using the four $\lambda$ values and the hyperparameters in Table II. This model that we refer to as Teacher is then used to teach two different student models with 160 and 128 filters per layer that we will refer to as Student-160 and Student-128, respectively.
Next, the teacher and the students are evaluated on the Kodak dataset following the same procedure described in related works like [20].
@@ IV-A Training the teacher and distilling the students
TABLE III: Number of GOPS required to process a 720p input after structured pruning
@@
Table III shows the number of floating point operations at inference time for each module in Fig. 2. Dropping about 16% of the teacher filters in Student-160 is enough to cut the computational complexity by a twofold factor; dropping a further 16% filters in Student-128 slashes the complexity to a fourth overall. The students are then fine-tuned as described in Sec. III-A to improve their RD performance. The teacher and the two distilled students are then pruned as in Sec. III-C, the parameters are quantized as in Sec. III-D, the GDN is implemented as in Sec. III-B, and the corresponding xmodels are compiled and individually ported to the ZCU102 FPGA for benchmarking.
IV-B Performance evaluation
Figure 9: Full R-D comparison with state of the art, [20, 22] and standard codecs on Kodak dataset
Coming to the comparison between FPGA-based models, we take as a reference the state of the art Sun2024 [22]. While in the hardware LIC literature it is not unseen comparing codecs measuring the PSNR delta for some target bitrate, we will always compare pairs of RD curves through Bjontegaard metrics[35]. Our Teacher outperform Sun2024 in RD efficiency (-8.95% BD-rate savings and 0.48 BD-PSNR better image quality), yet it is far from the 25 FPS target latency and below the 37 FPS scored by Sun-2024. Student-160 achieves however better RD efficiency than Sun2024 (-5.32 BD-Rate savings and 0.25% better image quality), and also better latency (43 vs 37 FPS). Student-128 is in the chart because while it is less RD efficient than Sun2024 (+7.80% BD-Rate savings and -0.10% image quality), it meets the 50 FPS bar and is very close to the 60 FPS bar too.
Fig.10 shows the corresponding RD curves on the CLIC dataset [32] of the main 192K quantized and pruned hardware model compared to the corresponding float model and the original reference model by Ballé. The drop in RD from float to hardware is to be expected and is consistent throughout the experiments we have observed, pruning and fixed point quantization result in model degradation, this drop is minimized using fine tuning on both tasks.
Table. IV gives the encoding results comparison with the literature, our proposed implementation yields the highest frame rate results at 43 FPS compared to the 37 FPS stare of the art, we achieve the fastest hardware level encoding using a smaller FPGA board and DSP slices, in turn this results in our proposed implementation having the lowest energy/frame consumption at 2.88 F/frame. Despite these advancements, it is worth noting that [22] achieves a higher DSP utilization efficiency of 91% compared to our 85%. This difference can be attributed to their exhaustive design space exploration, which optimize DSP allocation across the pipeline. Nonetheless, the trade-off in our design prioritizes overall throughput, RD performance and energy efficiency , achieving a better balance for real-time encoding scenarios. These numbers show how our approach outperforms state-of-the-art competitors like [22] both in RD-efficiency and latency terms and without the need for a complex hardware design space search.
TABLE IV: Hardware System Level Comparisons with the State of the Art
Clock (MHz) Encoding FPS ↑ DSP Usage DSP Efficiency ↑ Energy (J/frame) ↓
ZCU102 (ours) 300 43.6 2148 85% 2.88
ZCU104 [20] 300 3.90 NA NA 3.51
KU115 [21] 200 30.28 3706 54% 2.95
KU115 [22] NA 37.74 4160 91.7% NA
IV-C Analytical Model Validation
@@
  • Pipeline efficiency factor ($\eta$): 0.8 (Coarsely accounting for overhead of data transfer)
@@
For each model, we compute the FPS based on the workload and the given parameters.
Model 192K (528.2 GOP/frame)
$\text{FPS}=\frac{12,288\times 0.8}{528.2}\approx 18.6$
Model 160K (220.2 GOP/frame)
$\text{FPS}=\frac{12,288\times 0.8}{220.2}\approx 44.6$
Model 128K (153.0 GOP/frame)
$\text{FPS}=\frac{12,288\times 0.8}{153.0}\approx 64.1$
The FLOPs for each model are computed using the equation for $\text{FLOPs}_{\text{model}}$ based on the number of channels. The memory load is calculated using the memory traffic equation, and the FPS is estimated using the formulas described in earlier. The final FPS values are summarized in Table V.
@@ IV-C1 Discussion
The analytical model provides a reasonable estimate of FPS across configurations, with deviations of around 12%. It effectively captures the transition from compute-bound performance at higher $N$ values to memory-bound performance at lower $N$. For instance, at higher $N$ values, the performance is primarily determined by the computational resources available in the DPU cores, while at lower $N$ values, memory bandwidth plays a larger role. The framework allows users to select the number of channels ($N$) to meet rough FPS targets. For example, to achieve an FPS target of approximately 40, the model suggests $N=160$, balancing both computational and memory demands. For FPS greater than 60, reducing $N$ to 128 or less can minimize computational and memory requirements while still meeting throughput goals. However, we can see that for lower-load models like the 128K configuration, there is a greater discrepancy between the predicted and observed FPS. This is likely due to the fact that at lower $N$, the model approaches its memory limit, which hinders performance but it not accurately accounted for in the analytical model. Smaller models might be expected to improve throughput by utilizing more available resources, but in reality, performance is constrained by memory bandwidth at lower model sizes, resulting in memory-bound performance. Despite sufficient computational resources, the model cannot fully exploit them due to memory bandwidth limitations. Thus, while the analytical model suggests that reducing model size (lower $N$) alleviates computational load, this is not always the case when memory bandwidth becomes the limiting factor, leading to discrepancies, especially in low-computation configurations. The proposed analytical model offers a useful FPS estimate but simplifies key factors like memory access patterns, cache behavior, and intra-layer parallelism. It also doesn’t fully account for the dynamic interaction between memory bandwidth and compute resources or factors like resource contention and synchronization delays. These aspects require detailed, cycle-level analysis similar to [15] which we leave for future work.

### v5

@@ III-B Efficient GDN design for FPGA implementation
Other works [26, 27, 28, 29] replace the GDN layer with ReLU activation to simplify the implementation. For our FPGA-based LIC architecture, we are the first to propose a custom Generalized Divisive Normalization (GDN) hardware engine to optimize RD performance. The GDN layer is essential for reducing statistical dependencies between activations in LIC models, which improves compression quality. Given an input feature map $\mathbf{x}\in\mathbb{R}^{H\times W\times C}$, where $H$, $W$, and $C$ are the height, width, and number of channels respectively, the GDN operation is defined as:
@@ III-D Mixed precision quantization
The parameters that survived pruning are then quantized towards integers arithmetic operations over FPGA. In most hardware LIC implementations [27, 29, 26], all parameters are quantized to 8-bit precision to minimize latency. Conversely, we propose a mixed-precision quantization scheme where different model components are assigned different bit numbers. Specifically, the Generalized Divisive Normalization (GDN) layer is assigned a higher precision of 32-bit, while the rest of the model operates at 8-bit precision. This decision is motivated by the fact that the GDN layer, while only responsible for approximately 4% of the total FLOPS in the model, plays a critical role in improving the RD efficiency [36]. We recall that GDN’s primary function is to normalize activations across channels, promoting decorrelation in the latent space and improving RD efficiency. Lower precision (8-bit) introduces quantization noise in GDN’s operations such as division and normalization, which diminishes the accuracy of these transformations. By assigning 32-bit precision to the GDN layer, we reduce this quantization noise, leading to better normalization and decorrelation, ultimately enhancing rate-distortion performance, as we experimentally show also later on. Fig.6 illustrates the overall quantization strategy. Preliminarily, min and max values are extracted from the different model layers, allowing the quantizer to derive the optimal scale and zero point from the float model using the calibration subset. This subset includes 1000 random images from CLIC [37], DIV2K [38], and Open Images [39], used only for a forward pass. Symmetric quantization is achieved by setting the zero point to zero, with saturation applied to minimize quantization differences. As quantization impairs the model performance, the model undergoes a Quantization Aware Tuning (QAT) is then tuned again for 1000 iterations over a subset of 5000 images from CLIC, DIV2K, and Open Images. Simulated quantization operations are introduced on both weights and activations, except for the GDN layer, which maintains 32-bit precision to avoid losing the benefits of accurate normalization. During each forward pass, operations use the quantized weights and activations, and gradients are computed with respect to the quantized parameters using a Straight-Through Estimator (STE) [40] to approximate gradients through the quantization process. The outcome of the quantization process is an LIC model that exclusively relies on integer operations, meeting the requirements for efficient FPGA deployment.
@@ III-E Design Space Exploration
In this section, we finally describe the hardware configuration we use for implementing the quantized LIC model model on FPGA using the Vitis AI platform [34]. Namely, we target the Xilinx ZCU102 FPGA that houses three DPU cores running at 300 MHz. We propose below a fully pipelined configuration employing the common B4096-DPU core configuration. Such configuration provides a peak of 4096 Multiply-Accumulate (MAC) operations per cycle featuring 16 input channels, 16 output channels, and processing 8 pixels per cycle. Namely, we seek a fully pipelined configuration where each of the three DPU cores operates in parallel, as depicted in Fig.7. By dividing the image into 256 × 256 patches and processing multiple patches concurrently, the fully pipelined configuration optimizes resource utilization and minimizes end-to-end latency. Each DPU core is assigned a distinct patch, ensuring continuous processing and efficient use of computational resources.
This setup dedicates each DPU to processing different image patches simultaneously, maximizing throughput and reducing idle times. In contrast, the patch-based pipeline in [26] does not fully utilize available resources by only using the cores sequentially, leading to significantly lower throughput. Unlike [29, 28], we do not perform extensive design space exploration beyond efficient pipelining. Instead, we focus on reducing the computational load of the models, enabling us to outperform the works that rely on heavy design space exploration—without without requiring extensive hardware expertise or specialized knowledge.
@@
The model is compiled into target-specific DPU instructions in the form of an xmodel. The XIR compiler constructs an internal computation graph as an intermediate representation (IR), it then performs multiple optimizations, such as computation node fusion, and efficient instruction scheduling by exploiting inherent parallelism or exploiting data reuse on the specific target. This unified IR is broken down into subgraphs based on the corresponding control flow, for our models multiple DPU subgraphs are compiled as the custom GDN layers require separate graphs. Finally, an instruction stream is generated for each subgraph and everything is serialized into a target-specific, compiled xmodel file. This xmodel is encrypted and then deployed on the FPGA using the runtime API through the embedded processor. We deploy the models on the Zynq-UltraScale+ MPSoC DPU architecture -DPUCZDX8L, seen Fig.8. The nomenclature reflects the target FPGA, quantization bit-width, and design objective, where ’L’ denotes low latency. The DPU provides multiple configurations to trade off PL resources for performance. The B4096 core configuration yields the highest throughput with a peak of 18×16×16×2=4096 MAC/cycle. Three B4096-DPUCZDX8L cores are deployed on the ZCU102.
@@ III-F Analytical Model for Estimating FPS
$T_{\text{compute}}=\frac{N_{\text{workload}}}{\text{Peak Ops per Cycle}\times\eta}$ (5)
@@
The Peak Ops per Cycle is determined by the DPU configuration, which includes the number of input and output channels per layer, pixel parallelism, and core utilization. For each convolutional layer, the total number of operations per cycle per core is given by:
@@
For the DPU configuration with $PP=8$, $ICP=16$, and $OCP=16$, the peak operations per cycle per core (denoted as $\text{Ops}_{\text{core}}$) is:
$\text{Ops}_{\text{core}}=8\times 16\times 16\times 2=4096$
With three DPU cores operating in parallel, the total peak operations per cycle (denoted as $\text{Ops}_{\text{total}}$) is:
$\text{Ops}_{\text{total}}=3\times 4096=12,288$
@@
$T_{\text{frame}}=\frac{N_{\text{workload}}}{\text{Peak Ops per Cycle (total)}\times\eta}$ (6)
@@
FPS=Peak Ops per Cycle ($\text{FPS}=\frac{\text{Peak Ops per Cycle (total)}\times\eta}{N_{\text{workload}}}$ (7)
@@ IV Experiments and Results
IV-A Training the Hardware Efficient Models
IV-A1 Knowledge Distillation
To begin, we train the LIC model [8] from scratch to validate our experimental setup, following the original paper’s methodology. Our implementation builds upon the CompressAI framework [41], which provides a standardized baseline for benchmarking. For transparency, we report the hardware specifications: training was conducted on an A100-PCIE-40GB GPU, using the hyperparameters presented in Table I, ensuring reproducibility and substantiating our claim of training from scratch. The model is trained using a randomized train/validate/test split of the combined CLIC, DIV2K, and a subset of the Open Images dataset. Training is performed over batches of 256 × 256 randomly cropped patches, using four $\lambda$ values and the hyperparameters listed in Table I.
We refer to this model as the Teacher, which is then used to train several student models with 160 and 128 filters per layer, denoted as Student-160 and Student-128, respectively, the knowledge distillation scheme hyperparameters are presented in Table I . Next, the teacher and the students are evaluated on the Kodak dataset following the same procedure described in related works like [26].
TABLE I: Hyperparameters for the KD Training Scheme
Hyperparameter Initial Fine Tuning
Training Early Late
Learning Rate 1e-3 1e-4 1e-6
Decay [Step, Rate] 0.5M , 0.9 25k , 0.9 25k , 0.9
Batch Size [256x256] 16 8 8
$\alpha$ NA 0.1 1.0
$\beta$ NA 1.0 0.1
$\gamma$ NA 0.5 0.5
Steps 5M 250K 250K
$\lambda$ 0.0016, 0.0032, 0.0075, 0.045
@@ IV-A1 Knowledge Distillation
Early in the fine-tuning process, we set $\alpha=1.0$ and $\beta=0.1$ to encourage the student latent space to match that of the teacher, thereby achieving a similar latent space entropy. As training progresses, $\alpha$ is reduced to 0.1 and $\beta$ is increased to 1.0 to promote visual quality in the output. The transition between these two phases is automated by monitoring performance plateaus; for example, once the latent space loss $\mathcal{L}_{\text{latent}}$ stabilizes, the weights are automatically adjusted to emphasize the output space. Empirically, setting each phase to 250Ksteps satisfies our requirements on the chosen dataset.
Throughout the distillation process, the weight $\gamma$ is maintained at 0.5 to balance the rate-distortion (RD) loss and prevent catastrophic forgetting of the original end-to-end compression task. Table I lists all the relevant hyperparameters for the distillation process. It is worth noting that the teacher’s pre-trained hyperprior model is reused throughout training.
Table II details the number of floating point operations (GFLOPs) required at inference time for each module depicted in Fig. 2. Notably, for the main model, the distillation process reduces the teacher’s filters by approximately 16% in Student-160, effectively halving the computational complexity; a further 16% reduction in Student-128 slashes the overall complexity by a factor of four.
The students are then fine-tuned as described in Sec. III-A to improve their RD performance. The teacher and the two distilled students are then pruned as in Sec. III-C, the parameters are quantized as in Sec. III-D, the GDN is implemented as in Sec. III-B, and the corresponding xmodels are compiled and individually ported to the ZCU102 FPGA for benchmarking.
TABLE II: Number of GFLOPs required to process a 720p input after structured pruning
@@
IV-A2 Pruning and Quantization
Once we have our smaller student models, we apply iterative channel pruning and mixed-precision quantization-aware training (QAT) as described in Sections III-C and III-D. The effects of pruning and quantization are presented in Tables III and IV, respectively.
Table III shows that channel pruning can reduce model size by approximately 30% and significantly decrease inference latency (by around 40%). It also cuts the FLOPS by roughly 30%, although this comes with a modest drop in rate-distortion (RD) performance (e.g., a 0.25 dB decrease in PSNR for Student-160).
Subsequently, we perform QAT for 1000 iterations using a subset of the original training data. As illustrated in Table IV, QAT further compresses the model (achieving around a 43% reduction in size for Student-160) and provides moderate improvements in inference latency, while the impact on RD performance remains minimal.
TABLE III: Effect of channel-wise pruning compared to the reference models
160 128 96
$\Delta$ Model size ↓ -33.2% -27.1% -27.1%
$\Delta$ Inference latency ↓ -39.0% -24.3% -29.4%
$\Delta$ Avg PSNR (dB) ↑ -0.25 -0.28 -0.41
TABLE IV: Effect of Quantization Aware Training (QAT) compared to the pruned float model
160 128 96
$\Delta$ Model size ↓ -42.73% -48.92% -49.01%
$\Delta$ Inference latency ↓ -12.63% -13.18% -14.05%
$\Delta$ Avg PSNR (dB) ↑ +0.108 -0.095 -0.108
IV-B Analytical Model Validation
@@
  • Pipeline efficiency factor ($\eta$): 0.8 (Coarsely accounting for overhead of data transfer and pipeline efficiency)
@@
For each model, we use equations 5, 6, 7 and 8 to compute the FPS based on the workload and the given parameters , the FLOPs for each model are computed using the equation for $\text{FLOPs}_{\text{model}}$ based on the number of channels. The memory load is calculated using the memory traffic equation, and the FPS is estimated using the formulas described in earlier. The final FPS values are summarized in Table V.
@@
The analytical model provides a reasonable estimate of FPS across different configurations, with deviations of around 12%. It effectively captures the transition from compute-bound performance at higher $N$ values to memory-bound performance at lower $N$ values. For instance, at higher $N$, performance is mainly determined by the computational resources available in the DPU cores, while at lower $N$, memory bandwidth becomes a more significant limiting factor. The framework allows users to select the number of channels ($N$) to meet specific FPS targets. For example, to achieve an FPS target of approximately 40, the model suggests $N=160$, balancing computational and memory demands. For FPS values greater than 60, reducing $N$ to 128 or lower can minimize both computational and memory requirements, while still meeting throughput goals. However, for lower-load models, such as the 128K configuration, there is a larger discrepancy between the predicted and observed FPS. This discrepancy likely arises because, at lower $N$, the model reaches its memory bandwidth limit, which is not fully accounted for in the analytical model. Although smaller models may theoretically improve throughput by utilizing more available computational resources, in practice, performance is constrained by memory bandwidth in such configurations. As a result, the model cannot fully exploit the available computational resources, leading to memory-bound performance. While the analytical model offers useful FPS estimates, it simplifies key factors such as memory access patterns, cache behavior, and intra-layer parallelism. It also does not fully consider the dynamic interaction between memory bandwidth and compute resources, or factors such as resource contention and synchronization delays. These aspects require more detailed, cycle-level analysis, similar to that in [20], which we leave for future work.
IV-C Performance evaluation
Figure 9: Full R-D comparison with state of the art, [26, 29, 28] and standard codecs on Kodak dataset
TABLE VI: Final FPGA, BD metrics, FPS, and system-level results of various compression methods and comparisons with related works
Method Model Data Type Performance FPGA
Frame Rate BD-PSNR BD-Rate
Balle Hyperprior Baseline [8] fp32 N/A N/A N/A N/A
JPEG NA fp32 N/A -5.74 dB +250.88% N/A
JPEG-2000 NA fp32 N/A -2.99 dB +112.60% N/A
Teacher Models
Teacher 192_SP [8] int8 19 FPS -1.33 dB +37.63% ZCU102
Teacher 192_MP [8] int8/int32 16 FPS -0.66 dB +17.62% ZCU102
Student Models
Student 160_SP [8] int8 47 FPS -1.68 dB +49.35% ZCU102
Student 128_SP [8] int8 64 FPS -1.94 dB +57.29% ZCU102
Student 160_MP [8] int8/int32 44 FPS -0.90 dB +24.53% ZCU102
Student 128_MP [8] int8/int32 59 FPS -1.23 dB +34.57% ZCU102
Related Works
Sun-2024 [29] [8] int8 38 FPS -1.16 dB +28.22% VCU118
Jia-2022 [26] [32] int8 4 FPS -0.51 dB +26.34% ZCU104
Chen-2024 [28] [7] int8 33 FPS -1.54 dB +37.12% ZCU102
Coming to the comparison between FPGA-based models, we take as a reference the state of the art results in [29]. While in the hardware LIC literature it is not unseen comparing codecs measuring the PSNR delta for some target bitrate, we will always compare pairs of RD curves through Bjontegaard metrics[42]. Our Teacher model outperforms [29] in rate-distortion (RD) efficiency, achieving -8.95% BD-rate savings and 0.48 dB BD-PSNR better image quality. However, due to its larger computational cost, it falls short of the 25 FPS target and remains below the 37 FPS achieved by [29].
In contrast, Student-160 improves both RD efficiency and latency. It achieves -5.32% BD-rate savings and 0.25 dB BD-PSNR improvement over [29], while also surpassing its speed (43 FPS vs. 37 FPS), demonstrating the effectiveness of our proposed knowledge distillation scheme.
Student-128, while less RD-efficient than [29] (+7.80% BD-rate and -0.10 dB BD-PSNR), meets the 50 FPS real-time threshold and is close to the 60 FPS target, making it a viable option for latency-critical applications.
On the other hand, [26] and [28] fail to meet the required standards for different reasons. [26] employs a sophisticated model based on [43], achieving strong RD performance but at the cost of extremely low speed (only 4 FPS). In contrast, [28] prioritizes speed, reaching 33 FPS, but relies on an outdated model [7], resulting in uncompetitive RD performance, for instance it under performs even our smallest student model.
Fig.10 shows the corresponding RD curves on the CLIC dataset [37] of the main 192K quantized and pruned hardware model compared to the corresponding float model and the original reference model by Ballé [8]. The drop in RD from float to hardware is to be expected and is consistent throughout the experiments we have observed, pruning and fixed point quantization result in model degradation, this drop is minimized using fine tuning on both tasks.
Table VII presents a comparison of encoding results with the literature. Our proposed implementation achieves the highest frame rate at 43 FPS, surpassing the state-of-the-art 37 FPS. We attain the fastest hardware-level encoding while using a smaller FPGA board and fewer DSP slices. As a result, our design also achieves the lowest energy consumption per frame at 2.88 J/frame, second only to [28], which does not provide details on the energy measurement process.
Despite these advancements, [29] achieves a higher DSP utilization efficiency of 91%, compared to our 85%. This difference stems from their exhaustive design space exploration, which optimizes DSP allocation across the pipeline and searched for optimal configurations. However, our approach prioritizes a balanced trade-off between throughput, RD performance, and energy efficiency, making it better suited for real-time encoding scenarios.
These numbers show how our approach outperforms state-of-the-art competitors like [29] both in RD-efficiency and latency terms and without the need for a complex hardware design space search.
TABLE VII: Hardware System Level Comparisons with the State of the Art
Model Clock (MHz) Encoding FPS ↑ DSP Usage DSP Efficiency ↑ Energy (J/frame) ↓
ZCU102 (ours) Mid Cost [8] 300 43.63 2148 85% 2.88
ZCU102 [28] Low Cost [7] 200 32.95 1522 78% 0.12
ZCU104 [26] High Cost [32] 300 3.90 NA NA 3.51
KU115 [27] Mid Cost [8] 200 25.04 2654 94% 2.95
VCU118 [29] Mid Cost [8] NA 37.82 5864 93.7% NA
KU115 [29] Low Cost [7] NA 43.2 3752 93.7% NA